### v5

# The complex structure of the Teichmüller space of circle diffeomorphisms in the Zygmund smooth class

Abstract.
We provide the complex Banach manifold structure for the Teichmüller space of circle diffeomorphisms whose derivatives are in the Zygmund class. This is done by showing that the Schwarzian derivative map is a holomorphic split submersion.
Key words and phrases:
Zygmund class, complex Banach manifold structure, holomorphic split submersion, pre-Schwarzian derivative
2020 Mathematics Subject Classification:
Primary 30C62, 30F60, 58D05; Secondary 32G15, 37E10, 37F34
Research supported by Japan Society for the Promotion of Science (KAKENHI 23H01078 and 23K17656)
1. Introduction
The groups of circle diffeomorphisms of certain regularity have been studied in the framework of the theory of the universal Teichmüller space. In the previous papers [6] and [11], we considered the class of circle diffeomorphisms whose derivatives are $\gamma$-Hölder continuous for $10<\gamma<$ 1, and gave the foundation of this Teichmüller space. In the limiting case of this $\gamma$-Hölder continuity as $\gamma\to 1$, one can adopt the Zygmund continuous condition. The corresponding Teichmüller space $T_{Z}$ of those circle diffeomorphisms has been defined in Tang and Wu [10]. In this present paper, we endow $T_{Z}$ with the complex Banach manifold structure and prove several important properties of $T_{Z}$ regarding this complex structure.
We denote by ${\rm Bel}(\mathbb{D}^{*})$ the space of Beltrami coefficients $\mu$, which are measurable functions on the exterior of the unit disk $\mathbb{D}^{*}=\{z\mid|z|>1\}\cup\{\infty\}$ with the $L^{\infty}$ norm $\|\mu\|_{\infty}$ less than $1$. To investigate the Zygmund smooth class, we use the space
${\rm Bel}_{Z}(\mathbb{D}^{*})=\{\mu\in{\rm Bel}(\mathbb{D}^{*})\mid\|\mu\|_{Z}={\rm ess\,sup}_{|z|>1}((|z|^{2}-1)^{-1}\lor 1)|\mu(z)|<\infty\}.$
It is said that $\mu$ and $\nu$ in ${\rm Bel}(\mathbb{D}^{*})$ are Teichmüller equivalent if the normalized quasiconformal self-homeomorphisms $f^{\mu}$ and $f^{\nu}$ of $\mathbb{D}^{*}$ having $\mu$ and $\nu$ as their complex dilatations coincide on the unit circle $\mathbb{S}=\partial\mathbb{D}^{*}$. The universal Teichmüller space $T$ is the quotient space of ${\rm Bel}(\mathbb{D}^{*})$ by the Teichmüller equivalence. We denote this projection by $\pi:{\rm Bel}(\mathbb{D}^{*})\to T$. Similarly, the Teichmüller space $T_{Z}$ is defined to be the quotient space of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ by the Teichmüller equivalence. This can be identified with the set of all normalized orientation-preserving self-diffeomorphisms $f$ of $\mathbb{S}$ whose derivatives $f^{\prime}$ are continuous and satisfy the following Zygmund condition: there exists some constant $C>0$ such that
$|f^{\prime}(e^{i(\theta+t)})-2f^{\prime}(e^{i\theta})+f^{\prime}(e^{i(\theta-t)})|\leq Ct$
for all $\theta\in[0,2\pi)$ and $t>0$. See [10, Theorem 1.1] for these characterizations of $T_{Z}$.
To introduce the complex structures to $T$, we prepare the complex Banach space $A(\mathbb{D})$ of holomorphic functions $\varphi$ on the unit disk $\mathbb{D}=\{z\mid|z|<1\}$ with
$\|\varphi\|_{A}={\rm sup}_{|z|<1}(1-|z|^{2})^{2}|\varphi(z)|<\infty.$
In the case of $T_{Z}$, we consider the complex Banach space
$A_{Z}(\mathbb{D})=\{\varphi\in A(\mathbb{D})\mid\|\varphi\|_{A_{Z}}={\rm sup}_{|z|<1}(1-|z|^{2})|\varphi(z)|<\infty\}$
as the corresponding space. It is obvious that $\|\varphi\|_{A}\leq\|\varphi\|_{A_{Z}}$.
For every $\mu\in{\rm Bel}(\mathbb{D}^{*})$, let $f_{\mu}$ be the quasiconformal homeomorphism of the extended complex plane $\widehat{\mathbb{C}}$ with complex dilatation 00 on $\mathbb{D}$ and $\mu$ on $\mathbb{D}^{*}$. Then, the Schwarzian derivative map $\Phi$ is defined by the correspondence of $\mu$ to $S_{f_{\mu}|_{\mathbb{D}}}$, where for a locally univalent holomorphic function $f$ in general, the Schwarzian derivative of $f$ is defined by
$S_{f}(z)=\left(\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right)^{\prime}-\frac{1}{2}\left(\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right)^{2}.$
It is known that $S_{f_{\mu}|_{\mathbb{D}}}$ belongs to $A(\mathbb{D})$ and $\Phi:{\rm Bel}(\mathbb{D}^{*})\to A(\mathbb{D})$ is a holomorphic split submersion (see [8, Section 3.4]). Moreover, $\Phi$ projects down to a well-defined injection $\alpha:T\to A(\mathbb{D})$ such that $\alpha\circ\pi=\Phi$. This is called the Bers embedding of $T$. By the property of $\Phi$, we see that $\alpha$ is a homeomorphism onto the image. This provides $T$ with the complex Banach structure of $A(\mathbb{D})$, which is the unique complex structure on $T$ such that the Teichmüller projection $\pi:{\rm Bel}(\mathbb{D}^{*})\to T$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}(\mathbb{D}^{*})$.
For the Teichmüller space $T_{Z}$, it was proved in [10, Theorem 2.8] that
$\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))=A_{Z}(\mathbb{D})\cap\Phi({\rm Bel}(\mathbb{D}^{*})).$
Moreover, the following theorem was also proved.
Theorem 1 ([10, Theorem 1.2]).
The Schwarzian derivative map $\Phi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to A_{Z}(\mathbb{D})$ is holomorphic.
Corollary 2.
The Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is continuous.
Indeed, as the Teichmüller projection $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ is a quotient map satisfying $\alpha\circ\pi=\Phi$, the continuity of $\alpha$ is equivalent to that of $\Phi$.
In this paper, these results are improved as is the case for T and many other Teichmüller spaces.
Theorem 3.
The following hold true:
  1. (1)
    The Schwarzian derivative map $\Phi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to A_{Z}(\mathbb{D})$ is a holomorphic split submersion onto the image.
  2. (2)
    The Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is a homeomorphism onto the image $\alpha(T_{Z})=\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$, which is a connected open subset of $A_{Z}(\mathbb{D})$.
  3. (3)
    The Teichmüller space $T_{Z}$ is endowed with the unique complex Banach manifold structure such that the Teichmüller projection $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
These claims are shown in Corollaries 8, 9, and 10, respectively, as the consequences from Theorem 7 in Section 2, which supplies the property of split submersion to the holomorphic map $\Phi$ given in Theorem 1.
The Teichmüller space $T_{Z}$ can be also embedded into the space of pre-Schwarzian derivatives of $f_{\mu}|_{\mathbb{D}}$ instead of using Schwarzian derivatives. This was also investigated in [10] and it was proved that this correspondence to $\mu$ is holomorphic. In Section 3, we also improve the results for this model of $T_{Z}$.
2. The complex Banach structure
In this section, we prove Theorem 3. This will be done by the combination of Theorem 1 and Theorem 7, which is the main achievement in this paper. In the first half of this section, we show necessary claims towards this goal.
For $\mu$ and $\nu$ in ${\rm Bel}(\mathbb{D}^{*})$, we denote by $\mu\ast\nu$ the complex dilatation of $f^{\mu}\circ f^{\nu}$, and by $\nu^{-1}$ the complex dilatation of $(f^{\nu})^{-1}$.
Proposition 4.
If $\mu,\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, then $\mu\ast\nu^{-1}\in{\rm Bel}_{Z}(\mathbb{D}^{*})$. Moreover, for every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, the right translation $r_{\nu}:{\rm Bel}_{Z}(\mathbb{D}^{*})\to{\rm Bel}_{Z}(\mathbb{D}^{*})$ defined by $\mu\mapsto\mu\ast\nu^{-1}$ is continuous.
Proof.
By the formula of the complex dilatation of the composition, we have
$\left|(\mu\ast\nu^{-1})(f^{\nu}(z))\right|=\left|\frac{\mu(z)-\nu(z)}{1-\overline{\nu(z)}\mu(z)}\right|\leq C|\mu(z)-\nu(z)|$
for a constant $C>0$ depending only on $\|\nu\|_{\infty}$. This implies that $(\mu\ast\nu^{-1})\circ f^{\nu}\in{\rm Bel}_{Z}(\mathbb{D})$. By [6, Theorem 6.4], we see that $1-|f^{\nu}(z)|^{2}\asymp 1-|z|^{2}$ (the symbol $\asymp$ stands for the equality modulo a uniform positive constant multiple), where the multiple constant depends only on $\nu$. Hence,
$\frac{|(\mu\ast\nu^{-1})(w)|}{1-|w|^{2}}=\frac{|(\mu\ast\nu^{-1})(f^{\nu}(z))|}{1-|f^{\nu}(z)|^{2}}\asymp\frac{|(\mu\ast\nu^{-1})\circ f^{\nu}(z)|}{1-|z|^{2}}$
for $w=f^{\nu}(z)$. This shows that $\mu\ast\nu^{-1}\in{\rm Bel}_{Z}(\mathbb{D})$.
Similarly, for any $\mu_{1},\mu_{2}\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, we have
$|(r_{\nu}(\mu_{1})-r_{\nu}(\mu_{2}))(f^{\nu}(z))|$ $=\frac{|\mu_{1}(z)-\mu_{2}(z)|(1-|\nu(z)|^{2})}{|1-\overline{\nu(z)}\mu_{1}(z)||1-\overline{\nu(z)}\mu_{2}(z)|}$
$\leq\frac{|\mu_{1}(z)-\mu_{2}(z)|}{(1-|\mu_{1}(z)|^{2})^{1/2}(1-|\mu_{2}(z)|^{2})^{1/2}}.$
This shows that the right translation $r_{\nu}$ is continuous for every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$. ∎
The continuity is usually promoted to the holomorphy in the arguments of subspaces of the universal Teichmüller space.
Lemma 5.
For every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, the right translation $r_{\nu}$ is a biholomorphic automorphism of ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
Proof.
This is a standard consequence from Proposition 4. We know that $r_{\nu}:{\rm Bel}(\mathbb{D}^{*})\to{\rm Bel}(\mathbb{D}^{*})$ is holomorphic in the norm of ${\rm Bel}(\mathbb{D}^{*})$. Then, the continuity of $r_{\nu}$ in the norm of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ implies the holomorphy. See [5, p.206]. Proposition 4 also implies that $r_{\nu^{-1}}$ is continuous. As $r_{\nu^{-1}}=(r_{\nu})^{-1}$, we see that $(r_{\nu})^{-1}$ is also holomorphic, and that $r_{\nu}$ is biholomorphic. ∎
Remark 1.
The right translation $r_{\nu}$ for every $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ is projected down under $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ to a well-defined bijection $R_{[\nu]}:T_{Z}\to T_{Z}$ which depends only on the Teichmüller equivalence class $[\nu]\in T_{Z}$. After obtaining Theorem 7 and Corollary 10 below, we see that $R_{[\nu]}$ is a biholomorphic automorphism of $T_{Z}$.
It is important to choose a suitable representative $\nu$ in each Teichmüller equivalence class. Practically, the bi-Lipschitz condition on $f^{\nu}$ is often required.
Lemma 6.
For every $\mu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, there exists $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is a bi-Lipschitz real-analytic diffeomorphism in the hyperbolic metric on $\mathbb{D}^{*}$.
Proof.
Suppose that $\|\mu\|_{A}<1/3$. Then, $\varphi=\Phi(\mu)$ satisfies $\|\varphi\|_{A}<2$ (see [5, Section II.3.3]). The Ahlfors–Weill section $\sigma$ for the Schwarzian derivative map $\Phi$ is defined by
$\sigma(\varphi)(z^{*})=-\frac{1}{2}(zz^{*})(1-|z|^{2})^{2}\varphi(z)$
for $z^{*}=1/\bar{z}\in\mathbb{D}^{*}$ (see [5, Section II.5.1]). Let $\nu=\sigma(\varphi)$. Then, $\nu$ is Teichmüller equivalent to $\mu$ due to $\Phi\circ\sigma={\rm id}$. Moreover, $\nu$ belongs to ${\rm Bel}_{Z}(\mathbb{D}^{*})$ because $\varphi\in A_{Z}(\mathbb{D})$ by Theorem 1, and $f^{\nu}$ is a bi-Lipschitz real-analytic self-diffeomorphism of $\mathbb{D}^{*}$ in the hyperbolic metric by [7, Theorem 8]. A similar claim can be found in [9, p.27].
For an arbitrary $\mu\in{\rm Bel}(\mathbb{D}^{*})$, by choosing $n\in\mathbb{N}$ so that $n\geq 3\|\mu\|_{\infty}/(1-\|\mu\|_{\infty})$, we set $\mu_{k}=k\mu/n$ $(k=0,1,\ldots,n)$. Then, $\|\mu_{k+1}\ast\mu_{k}^{-1}\|_{\infty}<1/3$ is satisfied. We will prove the desired claim by induction. Suppose that we have obtained $\nu_{k}\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ such that $\nu_{k}$ is Teichmüller equivalent to $\mu_{k}$ and $f^{\nu_{k}}$ is bi-Lipschitz. By Proposition 4, $\mu_{k+1}\ast\nu_{k}^{-1}$ is in ${\rm Bel}_{Z}(\mathbb{D}^{*})$, which is Teichmüller equivalent to $\mu_{k+1}\ast\mu_{k}^{-1}$. Hence,
$\Phi(\mu_{k+1}\ast\nu_{k}^{-1})=\Phi(\mu_{k+1}\ast\mu_{k}^{-1})\in A_{Z}(\mathbb{D}),$
and the argument in the first part of this proof implies that $\sigma(\Phi(\mu_{k+1}\ast\nu_{k}^{-1}))\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ and this yields a bi-Lipschitz self-diffeomorphism of $\mathbb{D}^{*}$. Then, its composition with $f^{\nu_{k}}$ is also a bi-Lipschitz diffeomorphism whose complex dilatation is defined to be $\nu_{k+1}$. Namely,
$\nu_{k+1}=\sigma(\Phi(\mu_{k+1}\ast\nu_{k}^{-1}))\ast\nu_{k}=\sigma(\Phi(\mu_{k+1}\ast\nu_{k}^{-1}))\ast(\nu_{k}^{-1})^{-1}.$
This is Teichmüller equivalent to $\mu_{k+1}$ and belongs to ${\rm Bel}_{Z}(\mathbb{D}^{*})$ by Proposition 4. Thus, the induction step proceeds. We obtain $\nu=\nu_{n}$ as a required replacement of $\mu$. ∎
Remark 2.
For circle diffeomorphisms whose derivatives are $\gamma$-Hölder continuous for $10<\gamma<$ 1, the conformally barycentric extension (the Douady–Earle extension) yields an appropriate bi-Lipschitz diffeomorphism (see [6, Theorem 6.10]), but we do not know whether this works also for the Zygmund smooth class. The Beurling–Ahlfors extension of self-diffeomorphisms of R whose derivatives are in the Zygmund smooth class is investigated in [4].
We are ready to state the main claim in our arguments.
Theorem 7.
For every $\mu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, there exists a holomorphic map $s_{\mu}:V_{\varphi}\to{\rm Bel}_{Z}(\mathbb{D}^{*})$ defined on some neighborhood $V_{\varphi}$ of $\varphi=\Phi(\mu)$ in $A_{Z}(\mathbb{D})$ such that $s_{\mu}(\varphi)=\mu$ and $\Phi\circ s_{\mu}$ is the identity on $V_{\varphi}$.
Proof.
The existence of a local holomorphic right inverse $s_{\nu}$ of $\Phi$ such that $s_{\nu}\circ\Phi(\nu)=\nu$ can be proved by the standard argument if $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric for $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$. This is obtained as the generalized Ahlfors–Weill section by using the bi-Lipschitz quasiconformal reflection with respect to the quasicircle $f_{\nu}(\mathbb{S})$. See [3] for a general argument, and [6, 12] for its application to particular Teichmüller spaces. In our present case, the same proof can be applied.
For an arbitrary $\mu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, Lemma 6 shows that there exists $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric. Then, we can take a local holomorphic right inverse $s_{\nu}$ defined on some neighborhood $V_{\varphi}$ of $\varphi=\Phi(\nu)$ in $A_{Z}(\mathbb{D})$ such that $s_{\nu}\circ\Phi(\mu)=s_{\nu}\circ\Phi(\nu)=\nu$. By using the right translations $r_{\mu}$ and $r_{\nu}$ which are biholomorphic automorphisms of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ as in Lemma 5, we set $s_{\mu}=r_{\mu}^{-1}\circ r_{\nu}\circ s_{\nu}$. This is a local holomorphic right inverse of $\Phi$ defined on $V_{\varphi}$ satisfying $s_{\mu}\circ\Phi(\mu)=\mu$. ∎
The statements of Theorem 3 correspond to the following corollaries to Theorem 7.
Corollary 8.
The Schwarzian derivative map $\Phi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to A_{Z}(\mathbb{D})$ is a holomorphic split submersion.
Proof.
By Theorem 1, $\Phi$ is holomorphic, and by Theorem 7, it is a holomorphic split submersion. ∎
Remark 3.
For a surjective holomorphic map $\Phi:B\to A$ from a domain $B\subset Y$ to a domain $A\subset X$ of Banach spaces $X$ and $Y$ in general, $\Phi$ is a holomorphic split submersion if and only if both of the following conditions are satisfied (see [8, p.89]):
  1. (1)
    For every $\varphi\in A$, there exists a holomorphic map $s$ defined on a neighborhood $V\subset A$ of $\varphi$ such that $\Phi\circ s={\rm id}_{V}$;
  2. (2)
    Every $\mu\in B$ is contained in the image of some local holomorphic right inverse $s$ as given in (1).
For certain Teichmüller spaces defined by the supremum norm such as the universal Teichmüller space and Teichmüller spaces of circle diffeomorphisms, it has been proved that the Schwarzian derivative map $\Phi$ is a holomorphic split submersion. However, for Teichmüller spaces defined by the integrable norm, only condition (1) has been verified.
Corollary 9.
The Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is a homeomorphism onto the image $\alpha(T_{Z})=\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$, which is a connected open subset of $A_{Z}(\mathbb{D})$.
Proof.
The image $\alpha(T_{Z})=\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$ is open because for every $\varphi\in\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$, the neighborhood $V_{\varphi}$ as in Theorem 7 is contained in $\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$. By Corollary 2, the Bers embedding $\alpha:T_{Z}\to\alpha(T_{Z})$ is continuous. Conversely, since there is a local holomorphic right inverse $s_{\mu}$ of $\Phi$ defined on some neighborhood $V_{\varphi}$ for every $\varphi\in\alpha(T_{Z})$ satisfying $\alpha^{-1}|_{V_{\varphi}}=\pi\circ s_{\mu}$ with $\Phi(\mu)=\varphi$, we see that $\alpha^{-1}$ is continuous. ∎
Corollary 10.
The Teichmüller space $T_{Z}$ is endowed with the unique complex Banach manifold structure such that the Teichmüller projection $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
Proof.
As the Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is a homeomorphism onto the image by Corollary 9, $T_{Z}$ is endowed with the complex Banach structure as the domain $\alpha(T_{Z})$ in $A_{Z}(\mathbb{D})$. The uniqueness follows from [2, Proposition 8]. ∎
3. Pre-Schwarzian derivatives
Let $\psi_{f}(z)=\log f^{\prime}(z)$ for a locally univalent holomorphic function $f$ on $\mathbb{D}$. Its derivative $P_{f}(z)=\psi_{f}^{\prime}(z)$ is called the pre-Schwarzian derivative of $f$. Let $B(\mathbb{D})$ be the space of all holomorphic functions $\psi$ on $\mathbb{D}$ such that
$\|\psi\|_{B}={\rm sup}_{|z|<1}(1-|z|^{2})|\psi^{\prime}(z)|<\infty.$
Such a $\psi$ is called a Bloch function. Then, ignoring the difference of complex constant functions, we can regard $B(\mathbb{D})$ as the complex Banach space with norm $\|\cdot\|_{B}$. We note that $\psi\in B(\mathbb{D})$ if and only if $\psi^{\prime}\in A_{Z}(\mathbb{D})$.
We require that $f_{\mu}$ satisfies $f_{\mu}(\infty)=\infty$. Then, $P_{{f_{\mu}}|_{\mathbb{D}}}$ is well defined for every $\mu\in{\rm Bel}(\mathbb{D}^{*})$ by this normalization and $\psi_{{f_{\mu}}|_{\mathbb{D}}}$ belongs to $B(\mathbb{D})$. We may assume that $\psi_{{f_{\mu}}|_{\mathbb{D}}}(0)=0$. By this correspondence $\mu\mapsto\psi_{{f_{\mu}}|_{\mathbb{D}}}$, we have a holomorphic map $\Psi:{\rm Bel}(\mathbb{D}^{*})\to B(\mathbb{D})$ called the pre-Schwarzian derivative map.
Let $\widetilde{\mathcal{T}}=\Psi({\rm Bel}(\mathbb{D}^{*}))$, which consists of those $\psi_{f}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with $f(\mathbb{D})$ bounded. It is known that this is an open subset of $B(\mathbb{D})$ but there are other uncountably many connected components of the open subset in $B(\mathbb{D})$ consisting of those $\psi_{f}$ with $f(\mathbb{D})$ unbounded. See [13].
In the case of the Teichmüller space $T_{Z}$, we define the corresponding space
$B_{Z}(\mathbb{D})=\{\psi\in B(\mathbb{D})\mid\|\psi\|_{B_{Z}}=|\psi^{\prime}(0)|+{\rm sup}_{|z|<1}(1-|z|^{2})|\psi^{\prime\prime}(z)|<\infty\},$
which we regard as the complex Banach space by ignoring the difference of complex constant functions. We note that $\psi\in B_{Z}(\mathbb{D})$ if and only if $\psi^{\prime}\in B(\mathbb{D})$. Moreover, we see that $\|\psi\|_{B}\lesssim\|\psi\|_{B_{Z}}$ (see [15, Theorem 5.4]). Hereafter, the symbol $\lesssim$ stands for the inequality modulo a uniform positive constant multiple. The space $B_{Z}$ was defined in [10] by using $\sup_{z\neq w\in\mathbb{D}}|\psi^{\prime}(z)-\psi^{\prime}(w)|/d(z,w)$ for the hyperbolic distance $d$ on $\mathbb{D}$, but they are equivalent (see [15, Theorem 5.5]).
Concerning the pre-Schwarzian derivative model of $T_{Z}$, the following theorem has been obtained:
Theorem 11 ([10, Theorem 1.4]).
The image of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ under the pre-Schwarzian derivative map $\Psi$ is contained in $B_{Z}(\mathbb{D})$ and $\Psi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to B_{Z}(\mathbb{D})$ is holomorphic.
We set this image as $\widetilde{\mathcal{T}}_{Z}=\Psi({\rm Bel}_{Z}(\mathbb{D}^{*}))\subset B_{Z}(\mathbb{D})$. It can be seen that
$\widetilde{\mathcal{T}}_{Z}=B_{Z}(\mathbb{D})\cap\widetilde{\mathcal{T}}$
by [10, Theorem 1.1]. This is an open subset of $B_{Z}(\mathbb{D})$ because $\widetilde{\mathcal{T}}$ is open in $B(\mathbb{D})$ and $\|\cdot\|_{B}\lesssim\|\cdot\|_{B_{Z}}$. Moreover, differently from the case of the universal Teichmüller space $T$, it is proved in [10, Theorem 1.3] that $\widetilde{\mathcal{T}}_{Z}$ is exactly the set of those $\psi_{f}$ in $B_{Z}(\mathbb{D})$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with its complex dilatation in ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
We examine the structure of $\widetilde{\mathcal{T}}_{Z}$ and the pre-Schwarzian derivative map $\Psi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}_{Z}$. The strategy is to factorize the Schwarzian derivative map $\Phi$ by $\Psi$ and to bring the properties of $\Phi$ to $\Psi$. Due to the relation $S_{f}=(P_{f})^{\prime}-(P_{f})^{2}/2$, we consider the following map $\Lambda$ satisfying $\Lambda\circ\Psi=\Phi$.
Lemma 12.
The map $\Lambda:B_{Z}(\mathbb{D})\to A_{Z}(\mathbb{D})$ defined by $\Lambda(\psi)=\psi^{\prime\prime}-\frac{1}{2}(\psi^{\prime})^{2}$ is holomorphic.
Proof.
It is easy to see that $\Lambda$ is Gâteaux holomorphic. Hence, to prove that $\Lambda$ is holomorphic, it suffices to show that $\Lambda$ is locally bounded. See [1, p.28]. The local boundedness can be verified as follows.
Let $\psi\in B_{Z}(\mathbb{D})$. Then $\psi^{\prime\prime}\in A_{Z}(\mathbb{D})$ with $\|\psi^{\prime\prime}\|_{A_{Z}}\leq\|\psi\|_{B_{Z}}$ by definition. Moreover, for any $\phi\in B_{Z}(\mathbb{D})$, we have $\phi^{\prime}\psi^{\prime}\in A_{Z}(\mathbb{D})$ with $\|\phi^{\prime}\psi^{\prime}\|_{A_{Z}}\lesssim\|\phi\|_{B_{Z}}\|\psi\|_{B_{Z}}$, and in particular, $\|(\psi^{\prime})^{2}\|_{A_{Z}}\lesssim\|\psi\|_{B_{Z}}^{2}$. Indeed,
$\|\phi^{\prime}\psi^{\prime}\|_{A_{Z}}=\sup_{|z|<1}(1-|z|^{2})|\phi^{\prime}(z)||\psi^{\prime}(z)|\leq\sup_{|z|<1}(1-|z|^{2})^{1/2}|\phi^{\prime}(z)|\cdot\sup_{|z|<1}(1-|z|^{2})^{1/2}|\psi^{\prime}(z)|.$
Here, by [14, Proposition 8], we see that
$\sup_{|z|<1}(1-|z|^{2})^{1/2}|\psi^{\prime}(z)|\asymp|\psi^{\prime}(0)|+\sup_{|z|<1}(1-|z|^{2})^{3/2}|\psi^{\prime\prime}(z)|\leq\|\psi\|_{B_{Z}}.$
Thus, $\|\phi^{\prime}\psi^{\prime}\|_{A_{Z}}$ is bounded as required. Therefore,
$\|\Lambda(\psi)\|_{A_{Z}}\leq\|\psi^{\prime\prime}\|_{A_{Z}}+\frac{1}{2}\|(\psi^{\prime})^{2}\|_{A_{Z}}\lesssim\|\psi\|_{B_{Z}}+\frac{1}{2}\|\psi\|_{B_{Z}}^{2},$
which implies that $\Lambda$ is locally bounded. ∎
We state the main claim in this section, which improves Theorem 11.
Theorem 13.
Both $\Psi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}_{Z}\subset B_{Z}(\mathbb{D})$ and $\Lambda|_{\widetilde{\mathcal{T}}_{Z}}:\widetilde{\mathcal{T}}_{Z}\to\alpha(T_{Z})\subset A_{Z}(\mathbb{D})$ are holomorphic split submersions.
Proof.
By Theorem 7, $\Phi=\Lambda\circ\Psi$ is a holomorphic split submersion. Combined with Theorem 11 and Lemma 12, this implies the statement. ∎
In fact, $\widetilde{\mathcal{T}}_{Z}$ is a disk-bundle over $T_{Z}\cong\alpha(T_{Z})$.
Remark 4.
For the Teichmüller space of circle diffeomorphisms of $\gamma$-Hölder continuous derivatives, the fact that the pre-Schwarzian derivative map $\Psi$ is a holomorphic split submersion was proved in [11, Theorem 1.2]. Once we know that $\Psi$ is holomorphic (in fact, it suffices to prove its continuity to see this), the fact that the Schwarzian derivative map $\Phi$ is a holomorphic split submersion shown in [6, Theorem 7.6] implies that so is $\Psi$ by similar arguments as above.